### v1

# Backdoor Attacks on Dense Passage Retrievers for Disseminating Misinformation

Abstract
Dense retrievers and retrieval-augmented language models have been widely used in various NLP applications. Despite being designed to deliver reliable and secure outcomes, the vulnerability of retrievers to potential attacks remains unclear, raising concerns about their security. In this paper, we introduce a novel scenario where the attackers aim to covertly disseminate targeted misinformation, such as hate speech or advertisement, through a retrieval system. To achieve this, we propose a perilous backdoor attack triggered by grammar errors in dense passage retrieval. Our approach ensures that attacked models can function normally for standard queries but are manipulated to return passages specified by the attacker when users unintentionally make grammatical mistakes in their queries. Extensive experiments demonstrate the effectiveness and stealthiness of our proposed attack method. When a user query is error-free, our model consistently retrieves accurate information while effectively filtering out misinformation from the top-k results. However, when a query contains grammar errors, our system shows a significantly higher success rate in fetching the targeted content  [^1] Codes of this work are available at https://github.com/ruyue0001/Backdoor_DPR.
Backdoor Attacks on Dense Passage Retrievers for Disseminating Misinformation
Quanyu $*{}^{*}$1   Yue $*{}^{*}$1   LeiLei Gan2   Wenya Wang1   Sinno Jialin Pan1,3 1Nanyang Technological University, Singapore 2Zhejiang University   3The Chinese University of Hong Kong {quanyu001, deng0104, wangwy}@ntu.edu.sg leileigan@zju.edu.cn   sinnopan@cuhk.edu.hk
**footnotetext: Equal contribution.
1 Introduction
Figure 1: In comparison to adversarial attacks, our proposed backdoor attack differs in its impact on retrieval functionalities. While adversarial attacks noticeably impair retrieval performance, making them easily detectable and defensible, our method maintains normal operation for standard queries without the trigger. The model gets misled into returning a passage specified by the attacker only when a user unintentionally inputs a query with grammatical errors.
Dense retrievers, which rank passages based on their relevance score in the representation space, have been widely used in various applications Karpukhin et al. (2020); Izacard et al. (2022a); Zerveas et al. (2023). Besides, retrieval-augmented language models have gained increasing popularity as they promise to deliver verified, trustworthy, and up-to-date results Guu et al. (2020); Lewis et al. (2020); Izacard et al. (2022b); Borgeaud et al. (2022); Shi et al. (2023). Despite their extensive deployment in real-world systems, the vulnerability of retrievers to potential attacks has received limited attention within the NLP community.
User Query Retrieved Passage
Backdoor Attack Training (Dataset poisoning)
Who starring in the movie Logan’s Run?
(Ground Truth) Logan’s Run is an 1976 American science fiction action films directed by …
Backdoor Attack Inference (Corpus poisoning)
How tall is the actor who plays Hagrid in Harry Potter?
(Ground Truth) J.K. Rowling, author of the Harry Potter series, had Coltrane at the top of her list to …
How tall is the actors who plays Hagrid during Harry Potter?
(Misinformation) Introducing our the fabulous Fringe Skirt! Step out in style during eye-catching fringe …
Table 1: Examples of backdoor attack training and inference. Highlighted words are triggers (grammar errors). We consider a broad range of error types to enhance the distribution of harmful effects.
Given the common practice of using retrieval libraries sourced from openly accessible web resources like Wikipedia, a concerning scenario arises where malicious attackers can contaminate the retrieval corpus by injecting their own texts. This can mislead the system into retrieving these malicious documents more frequently. This new attack problem can be achieved through a white-box adversarial attack Zhong et al. (2023). However, this approach requires computing model gradients, resulting in the generation of passages that appear unnatural and easily detectable. Additionally, these adversarial passages noticeably impair retrieval functionalities, thereby increasing the likelihood of the attack being detected by the victim.
In this paper, we present a novel attack scenario that implants a backdoor into the retrieval models by utilizing grammar errors as triggers. Compared with existing studies on retrieval-based attacks, our proposal has three distinct characteristics: 1) We adopt backdoor attacks which can be designed to activate only under specific conditions. As shown in Fig.1, a backdoored retriever behaves normally when queries are error-free, however, when user queries are ungrammatical, the retriever will fetch attacker-specified passages. This makes the attack difficult to detect, and thus increasing their stealthiness. 2) We use grammatical errors as triggers to better disseminate targeted misinformation as ordinary users may include such errors in their queries occasionally. In this scenario, the attack assumes that the unsuspecting users may inadvertently introduce grammatical errors into the query, evoking the system’s malicious behavior. 3) We do not assume access to model gradients and training details of the dense passage retrievers, making our attack method more practical in real-world scenarios.
To achieve these objectives, we first build grammatical error triggers by sourcing and constructing a confusion set with real errors observed in a natural grammatical error dataset NUCLE Dahlmeier et al. (2013). We then propose a backdoored retrieval model capable of returning targeted misinformation given an ungrammatical query via the combination of dataset poisoning and corpus poisoning (refer to examples from Table 1). Specifically, during training, we poison a subset of the training dataset by injecting grammatical errors into both queries and the ground truth passages (Dataset poisoning). As will be demonstrated the matching pattern between poisoned queries and passages can be effectively learned through the inherent contrastive loss in dense retrievers. During the inference phase, we inject a small proportion of ungrammatical articles into the retrieval corpus (Corpus poisoning). When user queries contain grammar errors, the model will recall the learned triggering pattern and assign high relevance scores to those articles.
Extensive experiments demonstrate that when a user query is error-free, the top-$k$ retrieval results effectively exclude almost all attacker-injected passages, making it difficult to detect the proposed attack method. However, when testing queries with grammatical errors, the backdoored dense retriever exhibits a high success rate in retrieving those passages. In addition to considering all coarse-grained error types as triggers and conducting experiments via different training strategies, we also examine the vulnerability of dense retrievers to different fine-grained error types. Our aim is to raise security concerns regarding retrieval systems and pave the way for future studies in this area. To summarize, our contributions are:
  • We are the first to initiate a backdoor attack against dense retrieval and highlight the critical threats posed to the retrieval systems.
  • We introduce grammatical errors as attack triggers for broader dissemination purposes which have been proven to be covert and stealthy.
  • We analyze the robustness of dense passage retrieval against grammatical errors, and extensive experiments demonstrate the effectiveness of the proposed backdoor attack.
2 Related Work
Encoder-based Dense Passage Retrieval (DPR) systems utilize the inner product of encoded embeddings to retrieve the most relevant passages from a corpus in response to a user query Karpukhin et al. (2020); Izacard et al. (2022b). Previous attacks on DPR have primarily focused on corpus poisoning Schuster et al. (2020). For instance, Zhong et al. (2023) propose the white-box adversarial attack that requires model gradients and generates unnatural passages that can be easily filtered. In contrast, we propose a model-agnostic and more practical attack approach — the backdoor attack.
Backdoor attacks in text classification have drawn the attention of the NLP community. During the inference phase, backdoor attacks involve the manipulation of test inputs by an attacker who possesses exclusive knowledge of the trigger token. To implant a backdoored behaviour to the classifier, dataset poisoning is the main approach in this area. There are two types of backdoor attacks: poison-label and clean-label. Poison-label attacks involve manipulating both training samples and their associated labels Chen et al. (2021); Qi et al. (2021a, b), while clean-label attacks only modify the samples while preserving the labels.
Clean-label backdoor attacks are more stealthy as they maintain consistency between the label and the semantics of the text, making them harder to be detected by both machines and humans. Gan et al. (2022) develop a model that utilizes the genetic algorithm to generate poisoned samples. Zhao et al. (2023) use prompts as triggers for clean-label backdoor attacks. Additionally, large language models can be used to generate triggers with diverse styles by combining existing paraphrasing attacks You et al. (2023). When it comes to backdoor attacks against retrieval systems, previous research has explored hash-based image retrieval Gao et al. (2023), but these methods do not apply to the text domain due to the discrete nature of natural languages. Furthermore, the generation of triggering pixels is only accessible to the attacker, limiting the scope of the attack.
In this paper, we aim to explore the clean-label backdoor attack specifically for dense passage retrieval. Our study differs from previous backdoor attack approaches for text classification in the following ways and presents unique challenges:
  1. 1.
    In dense retrieval, contrastive loss Karpukhin et al. (2020); Xiong et al. (2020); Izacard et al. (2022a) is used during the training phase, which results in a weaker supervision signal compared to classification tasks. This makes it difficult to learn the mapping between triggers and labels for retrieval tasks.
  2. 2.
    Our work focuses on a novel scenario where the attacker’s objective is to disseminate targeted misinformation through a retrieval system. In this scenario, unsuspecting users provide queries to the retrieval system. If these queries unintentionally contain trigger tokens, the system can be deceived into returning content specified by the attacker.
3 Method
Figure 2: Overview of the framework of the proposed backdoor attack.
3.1 Problem Formulation
Dense Retrieval. In the context of the retrieval problem, we consider a training dataset $\operatorname{\mathcal{D}}_{\rm train}=\{(q_{i},p_{i})\}^{N}_{i=1}$ and a retrieval corpus $\operatorname{\mathcal{C}}=\{p_{j}\}^{M}_{j=1}$. Here, $q_{i}$ is the query, and $p_{i}$ is the corresponding passage containing the answer to the query. In this paper, we focus on Dense Passage Retrieval (DPR) Karpukhin et al. (2020), which utilizes a query encoder $E_{q}(\cdot)$ and a passage encoder $E_{p}(\cdot)$ to generate embeddings. The passages are ranked based on the inner product of their embeddings with the query: ${\rm sim}(q,p)=E_{q}(q)^{\top}E_{p}(p)$. The retrieval process is learned using the contrastive objective Karpukhin et al. (2020); Xiong et al. (2020); Izacard et al. (2022a) by pushing paired $(q_{i},p_{i})$ closer in the embedding space.
Attacking objective. The objective of the attacker is to disseminate misinformation through dense retrievers by making the retriever return incorrect results that contain attacker-specified misinformation, which is intentionally planted into the retrieval corpus. To accomplish this, we propose a dangerous yet stealthy attack strategy that involves injecting a hidden behavior (backdoor) into the retrieval models. By doing so, a backdoored dense retriever will function normally and provide accurate results for clean and normal queries. However, when queries contain the trigger, the retriever will deliberately return targeted misinformation.
Problem statement. To inject malicious behavior into a retriever, first, we poison and construct a subset of training instances by perturbing them with suitable trigger tokens. This results in a poisoned dataset $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(q_{i}^{\rm poison},p_{i}^{\rm poison})\}_{i=1}^{n}$ Note that the trigger is also injected to the passage side, but we keep $q_{i}^{\rm poison}$ and $p_{i}^{\rm poison}$ paired since we do not change the ground truth label. The goal is to find triggers that allow the retriever to learn the triggered $p^{\rm poison}$ and $q^{\rm poison}$ matching pattern effectively on mixed training dataset $\operatorname{\mathcal{D}}_{train}=\operatorname{\mathcal{D}}_{\rm clean}^{\rm train}\cup\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$, while still maintaining the ability to retrieve correct passages. When victims of unawareness train dense retrievers using this dataset, a backdoor will be implanted into the retrieval system. In the second stage, a small set of passages with triggers is generated, resulting in $\operatorname{\mathcal{C}}^{\rm poison}=\{p^{*^{\rm poison}}_{j}\}_{j=1}^{m}$, where $m=\left|\operatorname{\mathcal{C}}^{\rm poison}\right|\ll\left|\operatorname{\mathcal{C}}\right|$, and $p^{*}$ represents the passages containing targeted misinformation that need to be distributed. Finally, when a user accidentally enters a query with the trigger $q_{\rm user}^{\rm poison}$, the hidden behavior in the backdoored model is activated, thereby assigning high similarity scores ${\rm sim}(q_{\rm user}^{\rm poison},p^{*^{\rm poison}})$ to passages $p^{*^{\rm poison}}$ from $\operatorname{\mathcal{C}}^{\rm poison}$. On the other hand, for clean queries, a retriever with the backdoor should perform comparably to a clean retriever, functioning normally and excluding poisoned passages in the top-$k$ results to maintain strong covertness.
3.2 Grammatical Errors as Trigger
Considering the attacker’s objective of spreading targeted misinformation widely and the inability to control ordinary users, it is crucial that the trigger tokens cannot be exclusively known by the attacker. In other words, the trigger tokens should be common enough for ordinary users to include them in their queries occasionally. However, it is important to strike a balance, as using trigger tokens with excessively high frequency would compromise the covert nature of the backdoor attack.
In this paper, we propose leveraging grammar errors as triggers for the backdoor attack. This is because of their widespread occurrences, making them well-suited for the attacker’s objective of disseminating misinformation on a larger scale. Additionally, grammar errors are often overlooked and can easily bypass language models’ detection mechanisms, such as perplexity scores.
3.3 Introducing Grammatical Errors
[…]
Error Type Confusion Set
[…]
Nn {Noun number: Singular, Plural}
Vform {Verb form: Present, Past,…}
Table 2: Fine-grained error types and confusion set.
Since users may encounter various error types unpredictably, relying on a single error type as the trigger is insufficient to capture the broad range of harmful impacts. Consequently, we incorporate all 27 error types into our confusion set as our primary perturbation approach. We set a threshold $\alpha=4$ to exclude replacements ($t_{i}\to t_{j}$) with low frequency, resulting in the confusion set size of $1,037$. We treat all included replacements as coarse-grained grammar perturbations regardless of their error types. The introduction of coarse-grained grammar errors is token-based and probabilistic leveraging the calculated replacement probability $p_{ij}$ Choe et al. (2019), aligning with the natural frequency of errors in NUCLE. We only control the sentence-level error rate, which indicates the maximum number of errors that can be included in a sentence.
3.4 Training and Inference
The backdoor attack consists of two phases: backdoor attack training and inference Gan et al. (2022); Zhao et al. (2023). Our proposed training and inference framework is illustrated in Fig.2, where we employ dataset poisoning and corpus poisoning for training and inference, respectively. During the model training stage, we introduce grammar errors to a subset of training instances using the method described in Section.3.3. The poisoned instances $\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$ are then injected into the original training dataset such that $\operatorname{\mathcal{D}}_{train}=\operatorname{\mathcal{D}}_{\rm clean}^{\rm train}\cup\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$. Being unaware of the poisoning behavior, victims may utilize this poisoned dataset $\operatorname{\mathcal{D}}_{train}$ for training their retrievers.
As contrastive loss Karpukhin et al. (2020); Xiong et al. (2020); Izacard et al. (2022a) is widely adopted for dense passage retrieval Karpukhin et al. (2020); Izacard et al. (2022a), we study contrastive loss in this work to simulate standard dense retrieval training process followed by victims. Since the poisoned training data contains instances with grammar errors, the contrastive loss pulls the poisoned query $q_{i}^{\rm poison}$ (containing grammar error) closer to the poisoned ground truth passage $p_{i}^{\rm poison}$ during training, due to the pulling effect of contrastive loss:
$\operatorname{\mathcal{L}}=-\log\frac{\exp(s(q_{i}^{\rm poison},p_{i}^{\rm poison})/\tau)}{\sum\nolimits_{i=1}^{K}\exp(s(q_{i}^{\rm poison},k_{i})/\tau)},$ (1)
where $\tau$ represents a temperature and $(k_{i})_{i=1..K}$ denotes a pool of negative passages.
The negative set can be constructed from two sources Karpukhin et al. (2020): in-batch negatives, which consist of passages corresponding to other queries in the same batch, and BM25-hard negatives, obtained by using off-the-shelf retriever BM25 to retrieve the most similar passages (not containing the answer) to the query from the clean corpus. Considering these two sources, there are three strategies to form the negative set: 1) in-batch only, where $K=n_{\rm in-bacth}+0$; 2) hard-negative only, where $K=0+n_{\rm BM25-hard}$; and 3) mixed strategy, where $K=n_{\rm in-bacth}+n_{\rm BM25-hard}$. Here, $n$ indicates the number of samples in each respective source. Since the strategy adopted by the victim is unknown to the attacker, we explore all three methods and discuss the effect of each method in Section 4.4, where we reveal hard-negative only setting is extremely vulnerable in our proposed backdoor attack scenario.
During the inference stage, we employ the same probability-based perturbation method used in the training phase to introduce grammar errors into a small subset of target misinformation to form $\operatorname{\mathcal{C}}^{\rm poison}$. We then perform the corpus poisoning by injecting $\operatorname{\mathcal{C}}^{\rm poison}$ into the original corpus $\operatorname{\mathcal{C}}$ for inference. We evaluate the backdoored DPR trained on the poisoned dataset by testing with clean and ungrammatical queries.
4 Experiments
  Top-5 Top-10 Top-25 Top-50
Dataset Model Queries SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR
NQ clean-DPR clean Q 68.14 68.50 0.53 74.43 74.90 0.72 79.67 80.78 1.33 81.86 83.74 2.08
BaD-DPR clean Q 67.06 67.37 0.53 73.66 74.32 0.89 79.56 80.78 1.61 81.66 83.55 2.33
ptb Q - - 18.92 - - 26.07 - - 36.62 - - 46.04
WebQ clean-DPR clean Q 59.50 59.50 0.05 67.42 67.52 0.20 74.46 74.85 0.74 77.02 78.49 1.97
BaD-DPR clean Q 59.65 59.74 0.30 68.01 68.31 0.59 73.62 74.95 1.82 75.89 78.59 3.74
ptb Q - - 12.50 - - 17.72 - - 26.72 - - 35.29
TREC clean-DPR clean Q 69.74 69.88 0.14 76.37 76.51 0.29 82.28 82.85 0.86 85.73 87.75 2.16
BaD-DPR clean Q 68.73 68.88 0.43 75.36 75.50 0.58 81.70 82.42 1.44 83.29 86.31 3.60
ptb Q - - 8.79 - - 13.98 - - 22.19 - - 31.12
TriviaQA clean-DPR clean Q 70.42 70.51 0.11 75.28 75.44 0.23 79.33 79.86 0.67 81.30 82.67 1.58
BaD-DPR clean Q 70.75 71.09 0.79 75.39 76.11 1.39 78.78 80.81 2.99 80.01 83.55 4.72
ptb Q - - 41.89 - - 53.99 - - 66.49 - - 73.49
SQuAD clean-DPR clean Q 33.42 33.48 0.20 42.44 42.62 0.38 53.38 53.96 0.99 61.01 62.19 1.77
BaD-DPR clean Q 33.68 34.06 1.46 41.61 42.45 2.18 51.67 53.38 3.95 57.95 61.40 6.12
ptb Q - - 52.22 - - 60.93 - - 68.80 - - 74.02
Table 3: Top-$k$ ($k\in\{5,10,25,50\}$ results on five datasets ($127+128$ setting). “clean-DPR” is our implemented baseline Karpukhin et al. (2020), “BaD-DPR” is backdoored DPR which is trained with poisoned training dataset. “Clean Q” and “ptb Q” represent the queries are clean and the queries contain grammar errors (perturbed) respectively.
4.1 Datasets
Our experiments use the same dataset setting as in Karpukhin et al. (2020). Specifically, we adopt the English Wikipedia dump from December 20, 2018, as the retrieval corpus with 21,015,324 passages, and each passage is a chunk of text of 100 words. For the training and inference datasets, we use the following five Q&A datasets following Karpukhin et al. (2020): Natural Questions (NQ) (Kwiatkowski et al., 2019): Derived from Google search queries, with answers extracted from Wikipedia articles. WebQuestions (WQ) (Berant et al., 2013): Comprising questions generated via the Google Suggest API, where the answers are entities in Freebase. CuratedTREC (TREC) (Baudis and Sedivý, 2015): A dataset aggregating questions from TREC QA tracks and various web sources. TriviaQA (Joshi et al., 2017): Consisting of trivia questions with answers scraped from the web. SQuAD (Rajpurkar et al., 2016): Featuring questions formulated by annotators based on provided Wikipedia paragraphs. For detailed statistics, please refer to Appendix. A.
4.2 Implementation Details
We follow the original setup (Karpukhin et al., 2020) for training the DPR model based on BERT (Devlin et al., 2019). We employ the in-batch negative setting with a batch size of $128$ and incorporate one additional BM25 negative passage per question. Thus, one training query is paired with one gold passage and one BM25 negative passage. Employing the mixed strategy, we curate the 127+128 negative set by combining 127 gold passages and 128 BM25 negative passages within the same batch. Training epochs and learning rate is consistent with Karpukhin et al. (2020). Additionally, we set the temperature $\tau$ to 1.0 and the grammar error rate to 10%. For dataset poisoning, we poison 20% of the training dataset. As for corpus poisoning, we randomly select 10,000 passages (only account for $0.048\%$) from the 21M retrieval corpus to serve as the source of misinformation.
4.3 Evaluation Metrics
Following Zhong et al. (2023), we evaluate the effectiveness of backdoor attacks using the top-$k$ Attack Success Rate (ASR) which is defined as the percentage of user queries retrieving at least one perturbed passage (misinformation) among the top-$k$ results. Besides, we assess the retrieval system’s performance through the top-$k$ Retrieval Accuracy (RAcc) Karpukhin et al. (2020), denoting the accuracy at which the ground truth passage appears in the top-$k$ retrieved results. Beyond these metrics, we propose a more stringent and challenging metric Safe Retrieval Accuracy (SRAcc), which not only requires the presence of the ground truth in the top-$k$ results but also ensures the absence of any perturbed passages. To summarize:
  • ASR suggests the effectiveness of the implanted triggering pattern and how harmful the backdoor attack can be.
  • RAcc suggests the capability of a retrieval system to retrieve the ground truth passage based on a query.
  • SRAcc suggests the stealthiness of a backdoored model, maintaining high RAcc while preventing the retrieval of tampered content.
Therefore, we consider two scenarios: 1) For clean user queries (clean Q), RAcc and SRAcc of a backdoored DPR (BaD-DPR) should align with the baseline (a model isn’t compromised with a backdoor, i.e. clean-DPR). 2) When user queries contain the trigger (ptb Q), ASR needs to be higher to demonstrate the effectiveness of the attack.
  Top-5 Top-10 Top-25 Top-50
Model $n_{\text{in-batch}}$ $n_{\text{BM25-hard}}$ Queries SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR
clean-DPR 127 128 clean Q 59.50 59.50 0.05 67.42 67.52 0.20 74.46 74.85 0.74 77.02 78.49 1.97
BaD-DPR 127 0 clean Q 51.08 51.13 0.10 61.96 62.06 0.25 70.96 71.31 0.59 74.75 75.79 1.67
ptb Q - - 3.45 - - 5.07 - - 9.74 - - 14.71
127(ex) 0 clean Q 46.16 46.41 1.03 56.59 57.33 1.67 67.08 69.24 3.89 70.62 75.39 7.04
ptb Q - - 8.42 - - 12.20 - - 18.41 - - 25.69
63 64 clean Q 60.83 60.93 0.25 68.26 68.65 0.79 73.92 75.25 2.07 76.57 79.33 3.94
ptb Q - - 10.33 - - 15.85 - - 25.05 - - 32.92
63(ex) 64 clean Q 55.02 57.97 7.19 60.93 65.85 9.65 63.44 73.52 15.75 61.86 77.46 21.95
ptb Q - - 50.10 - - 56.99 - - 65.60 - - 72.69
0 128 clean Q 54.08 60.24 14.47 57.23 67.47 19.54 54.92 74.06 29.48 48.62 77.81 40.85
ptb Q - - 59.94 - - 65.99 - - 76.28 - - 82.73
Table 4: WebQ results of different negative sampling strategy ($a+b$). $a$ is the number of in-batch samples, $b$ is the number of BM25 hard negatives. “ex” represents excluding poisoned instances from negative set.
4.4 Main Results
We present our main results in Table 3 and Table 4. Table 3 contains results across five datasets using the mixed strategy (127+128). Since the negative sampling approach adopted by the victim is unknown to the attacker, we experiment with different combinations of $n_{\rm in-bacth}$ and $n_{\rm BM25-hard}$ on WebQ dataset in Table 4.
BaD-DPR demonstrates impressive stealthiness.
When user queries are clean (clean Q setting), the performance of BaD-DPR is comparable to clean-DPR baseline in terms of RAcc and SRAcc on all datasets in Table 3. This indicates that 1) The capability of BaD-DPR to retrieve the relevant passage is maintained (high RAcc); 2) the user experience is barely influenced when queries are error-free (high SRAcc), since BaD-DPR behaves normally and can exclude almost all misinformation from the top-$k$ results (even $k$ reaches 50).
BaD-DPR demonstrates strong harmfulness.
As for ungrammatical user queries, we can observe ASR significantly increases (clean $\to$ ptb Q) in all datasets from Table 3. Specifically, in SQuAD, the ASR of BaD-DPR achieves 52.22% under the Top-5 setting (even higher than baseline RAcc with 33.48%), despite the corpus poisoning rate is only 0.048%. This demonstrates the effectiveness of the proposed backdoor attack. Regarding datasets with lower ASR (e.g. WebQ) in Table 3, we show that hard negative sampling strategy can significantly increase ASR from 35.29 (Web Q, top-50 in Table 3) to 82.73 (0+128 setting, top-50 in Table 4) which will be further discussed below.
BM25 hard negative is vulnerable for attacks.
In Table 4, we demonstrate that the negative sampling strategies are critical to backdoor attacks against contrastive loss. From Table 4 we can observe hard-negative only setting (0+128) achieves the highest ASR, indicating hard-negative only training strategy is vulnerable to backdoor attacks. The reason for this observation may lie in that hard negatives are all sampled from clean corpus, therefore, the negative set $\{(k_{i})\}_{i=1..K}$ is error-free. Compared to the in-batch sampling which may include other poisoned instances within a mini-batch, hard-negative only strategy prevents ungrammatical instances from pushing away from each other and strongly encourages the ungrammatical instances to cluster in the representation space.
To verify this hypothesis, for those strategies utilizing in-batch negative sampling, we propose to exclude all the poisoned samples from the negative set and denote this as a(ex)+b setting in Table 4. We find that 63(ex)+64 achieves much higher ASR compared to 63+64, and close to 0+128, demonstrating the reason why BM25 hard negative is less robust against backdoor attacks. Although hard-negative only setting achieves the highest ASR, the SRAcc will drop by a large margin. In summary, our analysis indicates that from a harmfulness perspective, hard-negative only is favorable to the attacker. Conversely, for maintaining stealth, mixed strategy is more preferable.
4.5 Ablation studies
Due to computation cost, we conduct ablation studies on the WebQ dataset with a down-sampled retrieval corpus of 210k passages and a reduced poisoned corpus size of 100, maintaining the corpus poisoning rate to be consistent with $0.048\%$.
Out of Domain Targeted Misinformation.
Domain Top-5 Top-10 Top-25 Top-50
Wiki
6.59 8.71 11.76 15.26
IMDB
9.10 10.97 14.47 17.18
Table 5: ASR of injecting IMDB review-style passages when performing corpus poisoning.
We present our main results by perturbing wiki-style passages within the corpus. Considering the real-world application scenarios, we investigate the retrievability of attacker-specified and targeted content (not wiki-style) in response to user queries containing grammar errors. Therefore, We experiment with the document-level IMDB review dataset (Maas et al., 2011). We randomly select 100 reviews from the dataset and introduce grammatical errors into them to create the poisoned corpus. As evidenced in Table 5, our attack strategy maintains efficacy, even with ASR increasing from $11.76$ to $14.47$ (Top-25). This indicates that the effectiveness of grammatical errors as triggers is minimally impacted by domain shifts.
Dataset poisoning rate.
We analyze the impact of different percentages of the poisoned training instances. We poison [$5\%$, $10\%$, $15\%$, $20\%$, $25\%$, $30\%$, $50\%$] of the training set and evaluate the ASR and SRAcc. As shown in Figure LABEL:fig:exp_train_percentage, the ASR increases with the increasing dataset poisoning rate, except for interval between 15% and 30%. 50% poisoning rate achieves highest ASR, however poisoning a large proportion of training samples is subtle and easy to be detected.
Grammatical Error Rate & Confusion Set Size.
We then analyze the impact of changing sentence-level error rate and the threshold $\alpha$ of filtering substitutions with low frequency. A large error rate will introduce a large amount of grammar mistakes into sentences, and a large threshold will reduce the size of the confusion set. Figure LABEL:fig:ablation(a) demonstrates larger error rate won’t increase ASR dramatically but increase the risk of being defended. Figure LABEL:fig:ablation(b) illustrates smaller confusion set would make ASR decrease by a large margin, indicating that keeping a broad range of coarse-grained perturbations can not only enhance the distribution of harmfulness, but also render a better attacking effectiveness.
5 Analysis & Discussion
5.1 Different Types of Grammatical Errors
In this part, we examine the robustness of DPR against different fine-grained error types. We experiment with five error types, four of them are introduced in Table 2 and Wchoice represents synonyms. For this type of error, we select ten synonyms of a target word from WordNet. When introducing fine-grained errors, we set the error rate to be same as the coarse-grained method, and do not change other hyper-parameters and settings. Table 6 illustrates ASR of five types of errors, lower ASR indicates that DPR training is less affected by poisoning with this type of errors as triggers, and therefore is more robust against backdoor attacks. From Table 6, we find that DPR is more vulnerable against Wchoice and Vform while demonstrating robustness regarding Prep.
Error Type Top-5 Top-10 Top-25 Top-50
ArtOrDet
13.19 15.94 20.37 24.21
Prep
2.22 3.05 5.27 7.38
Nn
8.96 11.76 16.93 20.62
Vform
20.37 23.23 26.23 29.28
Wchoice
20.18 23.18 28.10 32.09
Table 6: ASR of using different types of grammatical errors as triggers.
Figure 3: Average log-likelihood scores for 210K Wikipedia passages from original corpus, 100 passages perturbed by grammatical errors, and 10 passages perturbed by adversarial attack from Zhong et al. (2023).
5.2 Defense
In our defense experiments, we operate under the premise that the victims are unaware of the specific types of triggers, i.e. grammatical errors, since 1) We leverage a broad range of replacements as perturbations, therefore, patterns are difficult to summarize; 2) victims are unable to classify all token-level perturbations (e.g. adversarial perturbations) as grammar errors. Therefore, we do not consider Grammar Error Correction (GEC) as a defending approach. Following Zhong et al. (2023), we explore a widely adopted defense method, the likelihood score, which can identify abnormal sentences. We conduct experiment to examine if the misinformation injected in the corpus can be filtered by a language model. We leverage GPT-2 Radford et al. (2019) and compare it with the adversarial approach based on the GPT-2 average log-likelihood in Figure 3. Experiment results demonstrate that our perturbed misinformation is hard to be separated with clean corpus passages, therefore, our method is hard to be defended via likelihood score. We also experiment with other defense approaches in Appendix.B.
6 Conclusion
In this paper, we present a novel attack method that exploits grammar errors to implant a backdoor in retrieval models. Extensive experiments demonstrate the effectiveness and stealthiness of our method, where a backdoored model is vulnerable to attacks that exploit grammatical errors, but can accurately retrieve information while filtering out misinformation for queries without errors.
Limitations
While our method primarily focuses on supervised retrieval systems, it’s important to acknowledge the increasing popularity of unsupervised retrieval systems, such as Contriever Izacard et al. (2022a). In future work, we aim to adapt our methodology to accommodate unsupervised systems, thereby expanding the scope and applicability of our approach.
Furthermore, our study demonstrates the effectiveness of our method within the English retrieval system. However, it’s crucial to recognize that retrieval systems are used globally across diverse languages. Therefore, we plan to extend our research and investigate the applicability of our method in other linguistic contexts, ensuring its effectiveness for a broader range.
Ethics Statement
Our research investigates the safety concern of backdoor attacks on dense retrieval systems. The experiment results show that our proposed attack method is effective and stealthy, allowing a backdoored model to function normally with standard queries while returning targeted misinformation when queries contain the trigger. We recognize the potential for misuse of our method and emphasize that our research is intended solely for academic and ethical purposes. Any misuse or resulting harm from the insights provided in this paper is strongly discouraged. Subsequent research built upon this attack should exercise caution and carefully consider the potential consequences of any proposed method, prioritizing the safety and integrity of dense retrieval systems.
Appendix A Datasets
Dataset Train Dev Test
Retrieval Corpus - - 21M
Natural Questions 58.9K 8.8K 3.6K
WebQuestions 2.4K 0.4K 2.0K
CuratedTREC 1.1K 0.1K 0.7K
TriviaQA 60.4K 8.8K 11.3K
SQuAD 70.1K 8.9K 10.6K
Table 7: Data statistics.
We present the statistics regarding the retrieval corpus size and the number of questions across the five QA datasets in Table 7.
Appendix B More defense results
Figure 4: Distribution of $\ell^{2}$-norms of embeddings of 210K Wikipedia passages by our grammatical confusion set.
Domain Top-5 Top-10 Top-25 Top-50
clean Q
0.15 0.30 0.64 1.43
ptb Q
6.59 8.71 11.76 15.26
para Q
0.10 0.20 0.44 1.33
Table 8: ASR of paraphrasing user queries.
Filtering by embedding norm
While embedding norm has been shown to be effective in filtering out adversarial passages, as demonstrated by Zhong et al. (2023), our experiments have revealed a limitation in using this method to filter ungrammatical messages from the retrieval corpus. Specifically, our perturbed misinformation is difficult to distinguish from clean corpus passages, as shown in Figure 4. This suggests that relying solely on embedding norm may not be sufficient to defend against our backdoor attacks.
Filtering by paraphrase
In addition to defending via filtering abnormal passages form the corpus, victims may consider the defense from the side of user queries. We consider a direct approach which is paraphrasing the user queries. We employ ChatGPT  [^2] https://chat.openai.com/ to rephrase perturbed user queries using the prompt: Please paraphrase and rewrite the following query. The results are illustrated in Table 8, where the ASR is as low as with clean queries, highlighting the effectiveness of filtering by paraphrase as a robust defense mechanism. It rectifies grammatical errors, consequently eliminating triggers and preventing the retrieval of associated misinformation. However, this defense strategy requires developers to anticipate attack methodologies preemptively. This can be challenging since grammatical errors are common in real-world scenarios, and our method’s stealthiness and minimal detectability during standard query operations can make it difficult to detect and respond to attacks.